OpenStudy (anonymous):

given f(x) = x^2 -x+1, find f(2+Δx)-f(2) / (Δx)

OpenStudy (anonymous):

i just need some help getting started on this one

OpenStudy (anonymous):

You need to replace in the variable x, with everything the f() says. So in your case, you would go like\[\frac{((2+ \Delta x)^2-(2+ \Delta x)+1)-(2^2-2+1)}{\Delta x}\]Now you start simplifying as much as you can :)

OpenStudy (anonymous):

ah ok. thanks!

OpenStudy (anonymous):

you're welcome, let me know if you need more help on this one.

OpenStudy (anonymous):

uh, i'm having a little trouble. i'm not sure if you noticed but i miswrote the original equation.

OpenStudy (anonymous):

Can you please provide the corrected equation then?

OpenStudy (anonymous):

it's in the open question area now

OpenStudy (anonymous):

Yes, I've used that equation that is currently on display to form the expression found in my answer. Do you need help simplifying it?

OpenStudy (anonymous):

hmmm, let me take another try at it. i'll post my results.

OpenStudy (anonymous):

alright then, remember that the first expression\[(2+ \Delta x)^2 \]is a polynomial of second degree so you need to expand that first.

OpenStudy (anonymous):

yes for that i got

OpenStudy (anonymous):

4+4(Δx)+(Δx)^2

OpenStudy (anonymous):

Yes, you're right. Now the second expression has a (-) sign infront so you need to make sure you factor that in,\[-(2+ \Delta x)\]What did you get for that?

OpenStudy (anonymous):

2+3(Δx)+(Δx)^2

OpenStudy (anonymous):

You mean the entire nominator looks like that? Like this?\[\frac{2+3 \Delta x + \Delta x^2}{\Delta x}\]

OpenStudy (anonymous):

oh i was just doing that one part

OpenStudy (anonymous):

You mean\[-(2+ \Delta x)=2 + 3 \Delta x+ \Delta x^2\]

OpenStudy (anonymous):

i tried doing it over and got confused. somewhere. i'm going to post what i have.

OpenStudy (anonymous):

ok let me see

OpenStudy (anonymous):

(4+4(Δx)+(Δx)^2)-2+Δx)+1)-(2^2-2+1) / Δx

OpenStudy (anonymous):

that's step 1

OpenStudy (anonymous):

correction

OpenStudy (anonymous):

(4+4(Δx)+(Δx)^2)-(2+Δx)+1)-(2^2-2+1) / Δx

OpenStudy (anonymous):

Ok you're good till here. From here, what did you do?

OpenStudy (anonymous):

then i have

OpenStudy (anonymous):

i missed a (

OpenStudy (anonymous):

(4+4(Δx)+(Δx)^2)-2-Δx+1)- 3 / Δx

OpenStudy (anonymous):

i simplified the right side. it made it seem easier for me.

OpenStudy (anonymous):

very good!

OpenStudy (anonymous):

Now, from here, how do you continue?

OpenStudy (anonymous):

i get 4(Δx)+(Δx)^2-2-Δx / Δx

OpenStudy (anonymous):

Ok but look, you have to separate the terms that look alike. For example, here's what I will do to simplify the numbers. I will look at the equation which looks like this\[\frac{((4+4 \Delta x+ \Delta x^2)-2-\Delta x+1)-3}{\Delta x}\]And i will pick out the numbers and I can see I have 4, (-2) and 1 and (-3). So I add them all up. Which gives me, 4-2+1-3=0 And this will remove all the numbers and leave out,\[\frac{(4 \Delta x+ \Delta x^2)-\Delta x}{\Delta x}\]Now remember, the 4 infront of Δx means we have 4 Δx-es. And in the equation there's another Δx that is being subtracted. What do you do?

OpenStudy (anonymous):

oh my gosh i totally messed up on the easy eliminating part. lol

OpenStudy (anonymous):

i forgot to take out the -2 lol

OpenStudy (anonymous):

Heh, no biggie. What's your progress to now?

OpenStudy (anonymous):

now i factor out Δx from the numerator and cancel it out with the denominator

OpenStudy (anonymous):

i get (Δx)(4+Δx-1) /Δx = 4+Δx-1

OpenStudy (anonymous):

opps

OpenStudy (anonymous):

Yes very good.. but are you forgeting something?

OpenStudy (anonymous):

i messed up. again lol.

OpenStudy (anonymous):

=3+Δx!!!

OpenStudy (anonymous):

Heheyy!! Great job qtexpress ;).. nice nickname lol

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

i just really had trouble figuring out how to start the equation. thanks for clearing that up!

OpenStudy (anonymous):

you're welcome, don't worry about that. it gets easier when you work with them for a while.

OpenStudy (anonymous):

can i show you my next question and you can tell me if i wrote the start of it correctly?